For a two tailed z-test using α = 0.05, the rejection region is defined as z values below or equal to ______ or z-values above or equal to _____.
Answer #1

solution:,

=0.05

/ 2 = 0.05 / 2 = 0.025

Z/2 = Z0.025 = +-1.96   ( Using z table )

defined as z values below or equal to __-1.96____ or z-values above or equal to __+1.96___